3. A rectangular (uncovered) container can fill up 3000 cm3 of liquid that fixed with a length of 20cm. Find the dimensions of the container that has the least surface area.
1
Expert's answer
2019-09-16T10:54:41-0400

Let length of the container = x = 20, width of the container = y, and height of the container = z.

Volume of the container "V=x\cdot y\cdot z=3000,\ y\cdot z=3000/20=150"

The surface of the container "S=x\cdot y+2(x+y)z=20y+2x\cdot z+2\cdot y\cdot z="

"20y+40\cdot150/y+300=20y+6000/y+300"

Using inequality of arithmetic and geometric means

"20y+6000/y\ge 2\sqrt{20y\cdot 6000/y}=400\sqrt{3}" , and equality possible when "20y=6000/y, y^2=300, y=\sqrt{300}" , therefor the area of the container is minimal when

"y=\sqrt{300}=10\sqrt{3},\ z=150/(10\sqrt{3})=5\sqrt{3}" .

Answer: dimensions of the container that has the least surface area are length = 20 cm, width = "10\sqrt{3}" cm, height = "5\sqrt{3}" cm.